# intersection.bib

### v2

@@ 5 Distance to find a road with large number of cars
In this section we analyse how far one needs to go from origin to see a road with large number of cars.
@@
Let $I$ denote the line segment with $\lfloor n^{1/3}\rfloor$ many vertices on $\mathcal{L}_{0}$ with midpoint $\boldsymbol{0}$. We divide the interval $(\epsilon,\frac{\pi}{2}-\epsilon)$ as we did in the proof of lower bound of Theorem 2.4. i.e., we fix large constants $M$ and $\ell_{0}$ which will be chosen later. For $1\leq i\leq[\frac{n^{1/3}}{2M}]$, we consider the intervals $A_{i}\subset(\epsilon,\frac{\pi}{2}-\epsilon)$, each of length $\frac{(\frac{\pi}{2}-2\epsilon)}{n^{1/3}}$ and midpoint $\frac{Mi(\frac{\pi}{2}-2\epsilon)}{n^{1/3}}$. Let us define the following events. For $1\leq i\leq[\frac{n^{1/3}}{2M}]$ we define
  • $\mathcal{K}_{i}:=\{\exists w\in I$ such that there are $\frac{n^{4/3}}{\ell_{0}}$ many vertices $v$ with $-2n\leq\phi(v)\leq-n,\theta_{v}\in A_{i}$ and $w\in\Gamma_{v}^{\theta_{v}}\}.$
@@
$\bigcup_{i=1}^{[\frac{n^{1/3}}{2M}]}\mathcal{K}_{i}\subset\{T_{n}\leq n^{1/3}\}.$
Same as before, considering the first two terms after applying inclusion-exclusion principle we have
$\mathbb{P}(T_{n}\leq n^{1/3})\geq\sum_{i=1}^{[\frac{n^{1/3}}{2M}]}\mathbb{P}(\mathcal{K}_{i})-\sum_{1\leq i<j\leq[\frac{n^{1/3}}{2M}]}\mathbb{P}(\mathcal{K}_{i}\cap\mathcal{K}_{j}).$ (5.1)
@@
But observe that, precisely this we have shown in the proof of Theorem 2.4 lower bound. Specifically, in (4.4) we have shown that on the event $\mathsf{TF}_{i}\cap\{E_{i}\leq\ell_{0}\}\cap\{Y_{n}-\frac{n^{4/3}}{4}\geq 0\}$ there is at least one vertex $w\in V_{M}$ such that there are $\frac{n^{4/3}}{4\ell_{0}}$ many $v\in\mathbb{Z}^{2}$ with $-2n\leq\phi(v)\leq-n,\theta_{v}\in A_{i}$ and $w\in\Gamma_{v}^{\theta_{v}}.$ So, we have
$\mathsf{TF}_{i}\cap\{E_{i}\leq\ell_{0}\}\cap\{Y_{n}-\frac{n^{4/3}}{4}\geq 0\}\subset\Bigl{\{}\sum_{1\leq j\leq\lfloor 4Mn^{1/3}\rfloor}\mathbbm{1}_{\mathcal{K}_{i}^{j}}\geq 1\Bigr{\}}.$
@@
So, we will work with $\mathcal{M}_{i}$ from now on. We will again apply an averaging argument in both space and time directions. But this time in the space direction we will average over disjoint intervals of length $n^{1/3}.$ Precisely, we do the following. We fix $m$ such that $-\lceil\frac{n}{100k_{i,j}}\rceil\leq m\leq\lceil\frac{n}{100k_{i,j}}\rceil$. We consider the line segments $\{v:|\psi(v)|\leq\lfloor n^{1/3}\rfloor\lfloor n^{1/3}\rfloor\}$ on $\mathcal{L}_{m}$ and partition them into line segments with $\lfloor n^{1/3}\rfloor$ many vertices. So, we get $\frac{c_{n}n^{4/3}}{100k_{i,j}}$ many mutually disjoint line segments each with $\lfloor n^{1/3}\rfloor$ many vertices for some uniformly bounded sequence $\{c_{n}\}$. We will denote the line segments by $I_{k,m}$, if the line segment lies on $\mathcal{L}_{m}$ and $k$ is the index to specify the partition in the space direction on $\mathcal{L}_{m}$. For all $-\lceil\frac{n}{100k_{i,j}}\rceil\leq m\leq\lceil\frac{n}{100k_{i,j}}\rceil$ and $1\leq k\leq 2\lfloor n^{1/3}\rfloor$ and for all $1\leq i\leq[\frac{n^{1/3}}{2M}]$ we define the following events.
@@
$\frac{c_{n}n^{4/3}}{100k_{i,j}}\mathbb{P}\left(\mathcal{M}_{i}\cap\mathcal{M}_{j}\right)=\sum_{\begin{subarray}{c}-\lceil\frac{n}{100k_{i,j}}\rceil\leq m\leq\lceil\frac{n}{100k_{i,j}}\rceil\\ 1\leq k\leq 2\lfloor n^{1/3}\rfloor\end{subarray}}\mathbb{P}\left(\mathcal{M}_{i}^{k,m}\cap\mathcal{M}_{j}^{k,m}\right)$
@@
Recall that we had define $N_{i,j}$ in the proof of lower bound in Theorem 2.5 and from (3.11) we have there exists constant $C>0$ such that for large $n$
$\frac{c_{n}n^{4/3}}{100k_{i,j}}\mathbb{P}\left(\mathcal{M}_{i}\cap\mathcal{M}_{j}\right)\leq\frac{Cn\log|k_{i,j}|}{k_{i,j}^{3}}.$
Combining all the above and from (5.1) we have there are $C,c>0$ (depending only on $\epsilon$)
$\mathbb{P}(T_{n}\leq n^{1/3})\geq\sum_{i=1}^{[\frac{n^{1/3}}{2M}]}\frac{c}{Mn^{1/3}}-\sum_{1\leq i<j\leq[\frac{n^{1/3}}{2M}]}\frac{C\log(M|i-j|)}{M^{2}|i-j|^{2}n^{1/3}}.$
@@
$\mathbb{P}(T_{n}\leq n^{1/3})\geq\frac{c}{M^{2}}-\frac{C}{M^{3}}.$
@@ 6.1.3 Figures
We provide Figures 8, 10 and 10 to illustrate the comparison between actual road segments and theoretical geodesic paths derived from our graph model. Each figure shows a map provided by ESRI World Topo Map [esriworldtopo2024] with a geodesic path, calculated based on elevation data and shortest path algorithms, overlaid on selected road segments. The intent behind these figures is to visually highlight where and how actual road networks align with or deviate from the calculated geodesic paths. This direct comparison serves as evidence for the hypothesis that road layouts tend to follow geodesic paths, considering elevation.
@@
One factor to consider is the graph model’s finite lattice structure, which simulates the continuous real-world terrain with a discrete grid. This approximation introduces potential errors in shortest path calculations, where the model’s sensitivity to edge weight value can manifest as notable deviations from expected paths. Such minor inaccuracies in terrain representation, therefore, might explain the observed misalignment between theoretical predictions and actual road configurations. This effect is made apparent between figures 10 and 10.
@@ 6.2.1 Calibrating the automated method
Figures 14, 14 and 14 refer to the startpoint near Cirencester (51.7076,-1.90897) in the South-East region. It is noted that sometimes the violin plot takes a peak when there are multiple traffic counts around the same distance. For example, consider the orange line in Figure 14. From the left side, it shows that there is a new maximum that is observed around 60 km after a peak at 40 km. But on the right side, traffic distribution shows that the traffic around 60 km is lower than the traffic at a distance of 40 km. This happens because multiple traffic counts of nearly the same values are observed near 40 km. The plots indicate that the appropriate strip width is between 3 km and 5 km.
Similarly, also in the South-East region Figures 17, 17 and 17 are from the startpoint (51.0826,-2.22697) near Gillingham. Figures 20, 20 and 20 are the plots with the startpoint near Wales (52.1577699,-3.5958771) in the sparser, West region of UK. Finally, Figure 21 is the plot with the startpoint near Glasgow, which belongs to the North region of UK. When taking the manual measurement for this plot, we get a high-traffic road within a distance of 10km, but when considering the automated measurement of a strip width of 10 km, we encounter another, much higher-traffic road, and these remain the busiest crossings until the distance of 150 km. This makes the runing maximum very different between the manual and the automated method. However, if we neglect the first few mismatched measurements, we get a good comparison. Figure 21 illustrates it all. As the points are randomly chosen, rarely bad points like this one are expected to occur, and we do not have an efficient way to eliminate this issue.
@@ 6.2.2 Statistical analysis
We were now ready to deploy the automated method in large scale. As mentioned above, the plots (Figures 14 - 21) indicate that the statistics are different between the South-East and the North and West regions of the UK. We therefore divide the UK into these two regions for our statistical analysis. Around 2000 random startpoints were considered inside the UK, and those points are at least 80 km horizontally away from big cities (like London, Birmingham, etc.), big rivers or oceans to avoid distorting effects to the measurements. For a fixed strip width and distance $d$, we observe how many points $x$ have traffic counts greater than $k$, and calculate the relative frequency by dividing by the total number of points, where $k$ runs from $1000$ to $100000$. Figures 23 and 23 show the traffic threshold encountered with relative frequency between $[0.49,0.51]$ as a function of the distance travelled on the Easterly line.
@@ Acknowledgements
The authors thank Riddhipratim Basu for useful discussions throughout the project. We also thank for discussions with Yuri Bakhtin, Atal Bhargava, Sanchari Goswami and Aquib Molla in the early phases of this work, and Janosch Ortmann later. MB was partially supported by the EPSRC EP/W032112/1 Standard Grant of the UK. SB was supported by scholarship from National Board for Higher Mathematics (NBHM) (ref no: 0203/13(32)/2021-R&D-II/13158). DH was partially supported by NSF grant DMS-2054559. KD was supported by the Prime Minister’s Research Fellowship PM/PMRF-22-18686.03 (PMRF ID: 0202965) from the Ministry of Education of India. This project was initiated at the International Centre for Theoretical Sciences (ICTS), Bengaluru, India during the program “First-passage percolation and related models” in July 2022 (code: ICTS/fpp-2022/7), the authors thank ICTS for the hospitality.

### v3

@@ 5 Distance to find a road with large number of cars
In this section we analyse how far one needs to go from origin to see a road with large number of cars. Recall that for a fixed constant $\ell_{0}$ we define
$T^{\ell_{0}}_{n}:=\min\left\{|\psi(v)|:v\in\mathcal{L}_{0},N_{v}\geq\frac{n^{4/3}}{\ell_{0}}\right\}.$
@@
Let $I$ denote the line segment with $\lfloor n^{1/3}\rfloor$ many vertices on $\mathcal{L}_{0}$ with midpoint $\boldsymbol{0}$. We divide the interval $(\varepsilon,\frac{\pi}{2}-\varepsilon)$ as we did in the proof of lower bound of Theorem 2.4. i.e., we fix large constants $M$ and $\ell_{0}$ which will be chosen later. For $1\leq i\leq[\frac{n^{1/3}}{2M}]$, we consider the intervals $A_{i}\subset(\varepsilon,\frac{\pi}{2}-\varepsilon)$, each of length $\frac{(\frac{\pi}{2}-2\varepsilon)}{n^{1/3}}$ and midpoint $\frac{Mi(\frac{\pi}{2}-2\varepsilon)}{n^{1/3}}$.
Let us define the following events. For $1\leq i\leq[\frac{n^{1/3}}{2M}]$ we define
  • $\mathcal{K}_{i}:=\{\exists w\in I$ such that there are $\frac{n^{4/3}}{\ell_{0}}$ many vertices $v$ with $-2n\leq\phi(v)\leq-n,\theta_{v}\in A_{i}$ and $w\in\Gamma_{v}^{\theta_{v}}$}.
@@
$\bigcup_{i=1}^{[\frac{n^{1/3}}{2M}]}\mathcal{K}_{i}\subset\{T^{\ell_{0}}_{n}\leq n^{1/3}\}.$
Same as before, considering the first two terms after applying the inclusion-exclusion principle (Proposition 6.1) we have
$\mathbb{P}(T^{\ell_{0}}_{n}\leq n^{1/3})\geq\sum_{i=1}^{[\frac{n^{1/3}}{2M}]}\mathbb{P}(\mathcal{K}_{i})-\sum_{1\leq i<j\leq[\frac{n^{1/3}}{2M}]}\mathbb{P}(\mathcal{K}_{i}\cap\mathcal{K}_{j}).$ (5.1)
@@
But observe that, precisely this we have shown in the proof of Theorem 2.4 lower bound. Specifically, in (4.4) we have shown that on the event $\mathsf{TF}_{i}\cap\left\{E_{i}\leq\frac{\ell_{0}}{4}\right\}\cap\{Y_{n}-\frac{n^{4/3}}{4}\geq 0\}$ there is at least one vertex $w\in V_{M}$ such that there are $\frac{n^{4/3}}{\ell_{0}}$ many $v\in\mathbb{Z}^{2}$ with $-2n\leq\phi(v)\leq-n,\theta_{v}\in A_{i}$ and $w\in\Gamma_{v}^{\theta_{v}}.$ So, we have
$\mathsf{TF}_{i}\cap\{E_{i}\leq\ell_{0}\}\cap\left\{Y_{n}-\frac{n^{4/3}}{4}\geq 0\right\}\subset\left\{\sum_{1\leq j\leq\lfloor 4Mn^{1/3}\rfloor}\mathbbm{1}_{\mathcal{K}_{i}^{j}}\geq 1\right\}.$
@@
So, we will work with $\mathcal{M}_{i}$ from now on. We will again apply an averaging argument in both space and time directions. But this time in the space direction we will average over disjoint intervals of length $n^{1/3}.$ Precisely, we fix $1\leq i<j\leq[\frac{n^{1/3}}{2M}]$. As before recall that $\theta_{i}$ and $\theta_{j}$ are midpoints of $A_{i}$ and $A_{j}$. Consider the parallelograms $W_{i,j}$ depending on $\theta_{i}$ and $\theta_{j}$ as defined in the proof of Theorem 2.5 lower bound with the smaller side length $n^{1/3}\cdot n^{1/3}$. We now fix $m$ such that $-\lceil\frac{n}{100k_{i,j}}\rceil\leq m\leq\lceil\frac{n}{100k_{i,j}}\rceil$ and consider $W_{i,j}\cap\mathcal{L}_{m}$. We consider these line segments of length $n^{1/3}\cdot n^{1/3}$ and partition them into line segments with $\lfloor n^{1/3}\rfloor$ many vertices. So, we get, for large $n$, at least $\frac{cn^{4/3}}{100k_{i,j}}$ many mutually disjoint line segments each with $\lfloor n^{1/3}\rfloor$ many vertices for some $c>0$ uniformly bounded away from 0. We will denote the line segments by $I_{k,m}$, if the line segment lies on $\mathcal{L}_{m}$ and $k$ is the index to specify the partition in the space direction on $\mathcal{L}_{m}$. For all $-\lceil\frac{n}{100k_{i,j}}\rceil\leq m\leq\lceil\frac{n}{100k_{i,j}}\rceil$ and $1\leq k\leq 2\lfloor n^{1/3}\rfloor$ and for all $1\leq i\leq[\frac{n^{1/3}}{2M}]$ we define the following events.
@@
$\frac{cn^{4/3}}{100k_{i,j}}\mathbb{P}\left(\mathcal{M}_{i}\cap\mathcal{M}_{j}\right)\leq\sum_{\begin{subarray}{c}-\lceil\frac{n}{100k_{i,j}}\rceil\leq m\leq\lceil\frac{n}{100k_{i,j}}\rceil\\ 1\leq k\leq 2\lfloor n^{1/3}\rfloor\end{subarray}}\mathbb{P}\left(\mathcal{M}_{i}^{k,m}\cap\mathcal{M}_{j}^{k,m}\right)$
@@
Recall that we had defined $N_{i,j}$ in the proof of lower bound in Theorem 2.5 and from (3.8) we have that there exists constant $C>0$ such that for large $n$
$\frac{cn^{4/3}}{100k_{i,j}}\mathbb{P}\left(\mathcal{M}_{i}\cap\mathcal{M}_{j}\right)\leq\frac{Cn\log|k_{i,j}|}{k_{i,j}^{3}}.$
Combining all the above and from (5.1) we have that there are $C,c>0$ (depending only on $\varepsilon$)
$\mathbb{P}(T^{\ell_{0}}_{n}\leq n^{1/3})\geq\sum_{i=1}^{[\frac{n^{1/3}}{2M}]}\frac{c}{Mn^{1/3}}-\sum_{1\leq i<j\leq[\frac{n^{1/3}}{2M}]}\frac{C\log(M|i-j|)}{M^{2}|i-j|^{2}n^{1/3}}.$
@@
$\mathbb{P}(T^{\ell_{0}}_{n}\leq n^{1/3})\geq\frac{c}{M^{2}}-\frac{C}{M^{3}}.$
@@ 6.1.3 Figures
We provide Figures 9, 11 and 11 to illustrate the comparison between actual road segments and theoretical geodesic paths derived from our graph model. Each figure shows a map provided by ESRI World Topo Map [esriworldtopo2024] with a geodesic path, calculated based on elevation data and shortest path algorithms, overlaid on selected road segments. The intent behind these figures is to visually highlight where and how actual road networks align with or deviate from the calculated geodesic paths. This direct comparison serves as evidence for the hypothesis that road layouts tend to follow geodesic paths, considering elevation.
@@
One factor to consider is the graph model’s finite lattice structure, which simulates the continuous real-world terrain with a discrete grid. This approximation introduces potential errors in shortest path calculations, where the model’s sensitivity to edge weight value can manifest as notable deviations from expected paths. Such minor inaccuracies in terrain representation, therefore, might explain the observed misalignment between theoretical predictions and actual road configurations. This effect is made apparent between Figures 11 and 11.
@@ 6.2.1 Calibrating the automated method
Figures 15, 15 and 15 refer to the startpoint near Cirencester (51.7076,-1.90897) in the South-East region. It is noted that sometimes the violin plot takes a peak when there are multiple traffic counts around the same distance. For example, consider the orange line in Figure 15. From the left side, it shows that there is a new maximum that is observed around 60 km after a peak at 40 km. But on the right side, traffic distribution shows that the traffic around 60 km is lower than the traffic at a distance of 40 km. This happens because multiple traffic counts of nearly the same values are observed near 40 km. The plots indicate that the appropriate strip width is between 3 km and 5 km.
Similarly, also in the South-East region Figures 18, 18 and 18 are from the startpoint (51.0826,-2.22697) near Gillingham. Figures 21, 21 and 21 are the plots with the startpoint near Wales (52.1577699,-3.5958771) in the sparser, West region of UK. Finally, Figure 22 is the plot with the startpoint near Glasgow, which belongs to the North region of UK. When taking the manual measurement for this plot, we get a high-traffic road within a distance of 10km, but when considering the automated measurement of a strip width of 10 km, we encounter another, much higher-traffic road, and these remain the busiest crossings until the distance of 150 km. This makes the runing maximum very different between the manual and the automated method. However, if we neglect the first few mismatched measurements, we get a good comparison. Figure 22 illustrates it all. As the points are randomly chosen, rarely bad points like this one are expected to occur, and we do not have an efficient way to eliminate this issue.
@@ 6.2.2 Statistical analysis
We were now ready to deploy the automated method in large scale. As mentioned above, the plots (Figures 15 - 22) indicate that the statistics are different between the South-East and the North and West regions of the UK. We therefore divide the UK into these two regions for our statistical analysis. Around 2000 random startpoints were considered inside the UK, and those points are at least 80 km horizontally away from big cities (like London, Birmingham, etc.), big rivers or oceans to avoid distorting effects to the measurements. For a fixed strip width and distance $d$, we observe how many points $x$ have traffic counts greater than $k$, and calculate the relative frequency by dividing by the total number of points, where $k$ runs from $1000$ to $100000$. Figures 24 and 24 show the traffic threshold encountered with relative frequency between $[0.49,0.51]$ as a function of the distance travelled on the Easterly line.
@@ Acknowledgements
The authors thank Riddhipratim Basu for useful discussions throughout the project. We also thank for discussions with Yuri Bakhtin, Atal Bhargava, Sanchari Goswami and Aquib Molla in the early phases of this work, and Janosch Ortmann later. We are also grateful for thorough reading of the manuscript and helpful comments on improvements by anonymous reviewers.
MB was partially supported by the Engineering and Physical Sciences Research Council EP/W032112/1 Standard Grant of the UK. SB was supported by scholarship from National Board for Higher Mathematics (NBHM) (ref no: 0203/13(32)/2021-R&D-II/13158). DH was partially supported by National Science Foundation grant DMS-2054559. KD was supported by the Prime Minister’s Research Fellowship PM/PMRF-22-18686.03 (PMRF ID: 0202965) from the Ministry of Education of India. This project was initiated at the International Centre for Theoretical Sciences (ICTS), Bengaluru, India during the program “First-passage percolation and related models” in July 2022 (code: ICTS/fpp-2022/7), the authors thank ICTS for the hospitality.
@@
Appendix: Auxiliary Estimates
Here we state the estimates that we have used in the proofs. For completeness, we first start with a simple special case of Bonferroni’s inequality.
Proposition 6.1.
Let $k\geq 2$ and $A_{1},\,A_{2},\,\dots A_{k}$ be any events in a probability space. Then
$\mathbb{P}\Bigl{(}\bigcup_{i=1}^{k}A_{i}\Bigr{)}\geq\sum_{i=1}^{k}\mathbb{P}(A_{i})-\sum_{1\leq i<j\leq k}\mathbb{P}(A_{i}\cap A_{j}).$
Proof.
We show this by induction. When $k=2$, the equality is the base case of the inclusion-exclusion formula. Let us now assume the statement for $k$, and write
$\mathbb{P}\Bigl{(}\bigcup_{i=1}^{k+1}A_{i}\Bigr{)}$ $=\mathbb{P}\biggl{(}\Bigl{(}\bigcup_{i=1}^{k}A_{i}\Bigr{)}\cup A_{k+1}\biggr{)}$
$=\mathbb{P}\Bigl{(}\bigcup_{i=1}^{k}A_{i}\Bigr{)}+\mathbb{P}(A_{k+1})-\mathbb{P}\biggl{(}\Bigl{(}\bigcup_{i=1}^{k}A_{i}\Bigr{)}\cap A_{k+1}\biggr{)}$
$=\mathbb{P}\Bigl{(}\bigcup_{i=1}^{k}A_{i}\Bigr{)}+\mathbb{P}(A_{k+1})-\mathbb{P}\Bigl{(}\bigcup_{i=1}^{k}(A_{i}\cap A_{k+1})\Bigr{)}.$
We can now apply the inductive hypothesis on the probability of first union, while Boole’s inequality on the second. With the minus sign of the display these point in the same direction:
$\mathbb{P}\Bigl{(}\bigcup_{i=1}^{k+1}A_{i}\Bigr{)}$ $\geq\sum_{i=1}^{k}\mathbb{P}(A_{i})-\sum_{1\leq i<j\leq k}\mathbb{P}(A_{i}\cap A_{j})+\mathbb{P}(A_{k+1})-\sum_{i=1}^{k}\mathbb{P}(A_{i}\cap A_{k+1})$
$=\sum_{i=1}^{k+1}\mathbb{P}(A_{i})-\sum_{1\leq i<j\leq{k+1}}\mathbb{P}(A_{i}\cap A_{j})$
as required by collapsing the summations. ∎
Next, the statement of Berry-Esseen inequality.
Proposition 6.2 (Berry-Essen inequality, [Feller_book, XVI.5.]).
Let $X_{1},X_{2},\dots$ be i.i.d. with $\mathbb{E}(X_{i})=0,\mathbb{E}X_{i}^{2}=\sigma^{2},\mathbb{E}|X_{i}|^{3}=\rho<\infty.$ If $F_{n}(x)$ is the distribution of $\frac{X_{1}+X_{2}+\dots+X_{n}}{\sigma\sqrt{n}}$ and $\mathcal{N}(x)$ is the standard normal distribution, then for all $n$ and $x$
$\left|F_{n}(x)-\mathcal{N}(x)\right|\leq\frac{3\rho}{\sigma^{3}\sqrt{n}}.$
Next we state the estimates from exponential LPP. We start with transversal fluctuation of semi-infinite geodesics.
Proposition 6.3.
[BBB23, Proposition 1.5] For $\varepsilon>0$ there exist $C_{1},c_{1}>0$ such that for all $T>0,\ell>0,n\geq 1$ and for all $\alpha\in(\varepsilon,\frac{\pi}{2}-\varepsilon)$ we have
  1. (i)
  2. (ii)
Next we state a proposition about the fact that geodesics tend to coalesce.
Proposition 6.4.
$\mathbb{P}(M_{n}^{k}\geq\ell)\leq Ce^{-c\ell^{1/128}}.$ (6.5)
Finally we have the following lemma about the size of intersection of geodesics with different directions. Recall the line segments $\mathcal{J}^{\prime}_{i},\mathcal{J}_{i}^{\prime\prime},\mathcal{J}_{j}^{\prime},\mathcal{J}_{j}^{\prime\prime}$ and the parallelogram $W_{i,j}$ in the proof of Theorem 2.5 lower bound. We have the following lemma.
Lemma 6.6.
[BBB23, Lemma 3.2, Lemma 3.3] Consider any geodesic $\Gamma_{i}$ (resp. $\Gamma_{j}$) starting from $\mathcal{J}_{i}^{\prime}$ (resp. $\mathcal{J}_{j}^{\prime}$) and ending at $\mathcal{J}_{i}^{\prime\prime}$ (resp. $\mathcal{J}_{j}^{\prime\prime})$. If $I_{\Gamma_{i},\Gamma_{j}}$ denote the intersection size of $\Gamma_{i}$ and $\Gamma_{j}$ inside $W_{i,j}$ then for $\ell$ and $n$ large enough there exist $C,c>0$ such that we have